# Graphical Abstract Highlights

### v1

@@ 3 Mathematical analysis of the model
The mathematical properties of the model are shown in Appendix (8).
4 Steady states and the reproductive number
The PDE system (7) admits a non-trivial disease-free equilibrium point such that
The uniqueness of can be verified using [(Wang et al., 2018), Lemma 2.1]. The basic reproductive number of the deterministic part of our model is computed following van den Driessche and Watmough (2002), as follows:
The matrix of new infections is given by while that of transitions is given by where
Taking partial derivatives and evaluating at the disease-free equilibrium point, we have
and
The basic reproductive number is given by the spectral radius of which yields
where
and
with equality when which shows the contribution of migratory birds in increasing the amplifying host population. Figure (2) shows the plot of the basic reproductive number all over Germany, as of 2024. shows a hotspot in the south-eastern region of Germany towards the northern part, which is in agreement with the observed data. It also identifies another hotspot in the south-west, around Bobenheim-Roxheim. The regions where is zero are regions where the temperature data has no values, mainly because of the presence of a sea. To validate the simulated we performed a chi-squared goodness of fit test at the 5% significance level. From this test, we obtained a p-value of which indicates that there are no significant differences between the observed and simulated patterns.
5 Simulation framework
The PDE model is simulated starting from 1 January 2019 until 5 December 2024 using a daily time step. Cases for the year 2018 (Figure 2(a)) are set as initial conditions at each location where cases were confirmed by 1 January 2019. To account for the inflow of new data through ongoing WNV surveillance activities, we model the spread every year, in such a way that the initial conditions for each year are the final solution of the previous year. This allows us to update the initial conditions in areas where the model did not predict cases, e.g. in 2019 a new case of WNV was confirmed in the north of Germany (Hamburg), and cases in this region may have played a role in the spreading pattern observed in the north of Germany during the subsequent years. In the process, we use inverse modeling to infer the diffusion and advection coefficients that could have contributed to the spreading pattern for each year. Temperature strongly influences the activity of Cx. pipiens, including their competence in transmitting WNV (Ciota et al., 2014), and thus, we included a temperature-dependent mosquito biting rate latency rate and mortality rate Hourly 2m temperature data from the ERA5 reanalysis is downloaded from the climate data store (Hersbach et al., 2023) and averaged to yearly data. During the simulation process, and are dynamically updated based on the temperature at different locations in Germany. To solve our PDE model, we use the Matlab PDEToolbox (The MathWorks, Inc., 2024), which uses a finite element algorithm. The toolbox solves PDEs of the form
and in our case, and matrices and are defined in Appendix (8.2). The matrix consists of the diffusion coefficients, while stores advection terms. The model is solved as follows (The MathWorks, Inc., 2024):
The map of Germany is downloaded as a shape file from (https://gadm.org/download_country_v3.html) and simplified using mapshaper (Harrower and Bloch, 2006). WNV infection data consisting of infected and dead birds and equids for the years 2018 - 2024 (Figure (2(a))) were obtained from the World Organisation for Animal Health (2018 - 2024) (WOAH). Human WNV cases data was obtained from European Centre for Disease Prevention and Control (2019, 2020, 2021, 2022, 2023, 2024). The observed data indicates presence data only, i.e., no absence data is available. For this reason, we qualitatively present our results (Low - High) to visualize the simulation results. To assess how well the model predictions match the observational data, we performed a chi-squared goodness of fit test at 5% level of significance, with the hypothesis:
There are no significant differences between the observed and simulated patterns.
There are significant differences between the observed and simulated patterns.
p-values are computed and the distributive patterns are tested for each year, while the null hypothesis is rejected whenever p-values are less than or equal to 0.05. The relationship between the model output and the observed data is further quantified using the cosine similarity index which measures the extent to which two vectors point in the same direction, without comparing their magnitudes (Huang et al., 2008). This index is computed using the formula where and represent the two datasets (modeled and observed) expressed as vectors. It is measured in the interval , where values closer to 1 indicate a high degree of similarity between the two vectors, while values close to zero indicate no similarity in the patterns described by these vectors. Values closer to -1 indicate that the two vectors point in opposite directions (Huang et al., 2008). Initial conditions and parameter values used in the simulations are summarized in Tables (1) and (2), respectively. We first simulate the model with migratory birds. It is indisputable that at every location with infectious resident and migratory birds or equids, infectious mosquitoes are also present. Therefore, we validated the model prediction using the distribution pattern of infectious resident birds
6 Results
The model initialized with 2018 WNV cases data shows strong visual agreement with observed patterns in Germany throughout the 2019 - 2024 simulation period, especially predicting hotspots in the eastern regions, including Berlin. The chi-squared test for 2019 indicated significant differences (p = 0.01), likely due to the model missing new cases in northern and central Germany; the cosine similarity was still moderately high at 0.6803 (Table 3). In subsequent years (2020 - 2024), p-values exceeded 0.05. Cosine similarity improved over the years, reaching 0.99 in 2024. Diffusion coefficients varied annually, showing a slight increase over the years, indicating higher mobility of both mosquito and host populations. Notably, in 2023 and 2024, the model accurately predicted cases in regions where none had been reported previously, hinting at its ability to anticipate under-reported spread patterns (Figure 4(e) - 4(f)). This capability was supported when 2024 observations confirmed some of the simulated 2023 circulation areas. The model’s performance suggests that incorporating spatial heterogeneity significantly enhances its predictive power. The diffusion speeds of both mosquitoes and hosts were positively correlated to temperature (Figure 5).
Year - squared statistical test Cosine similarity index Diffusion coefficient Value 0.0001 2019 <0.05 0.75 0.0003 0.0005 0.0002 2020 >0.05 0.68 0.00035 0.0004 0.00045 2021 >0.05 0.77 0.0005 0.00056 0.0005 2022 >0.05 0.73 0.00065 0.0007 0.05 2023 >0.05 0.80 0.06 0.07 0.05 2024 >0.05 0.99 0.065 0.07
6.1 Impact of migratory birds in disseminating WNV to distant regions:
During the years 2019 - 2024, the model is initially simulated with migratory birds. The advection vector is set to zero during the winter and summer seasons, as we assume that no migration is taking place. We then set the direction of flight to the northeast direction during spring, i.e., 1 March to 31 May of each year. During the autumn (1 September to 30 November), migratory birds start migrating towards the southwest, and we set the vector to indicate a southwest flow each year. Model results were the same for the cases when there were no migratory birds for all the years, except for the year 2023. During the year 2023, without migratory birds (Figure 6(a)), the model could not account for the new isolated cases observed in the western parts of Germany. After introducing migratory birds and their seasonal flow (Figure 6(b)), the model correctly predicts the observed pattern of WNV in 2023.
7 Discussions and conclusions
Understanding the WNV spreading pattern allows early detection of WNV outbreaks. The identification of hotspots also helps early planning and resource allocation for preventive measures. At the same time, it is expensive to carry out surveillance programs aiming to understand the spread of WNV. Therefore, in this study, we modeled the spread of WNV in Germany using a temperature-driven PDE model with advection and diffusion terms. In agreement with previous studies (Bhowmick et al., 2020; Bergsman et al., 2016; Mbaoma et al., 2024), at the temporal level, the basic reproductive number indicates that the presence of migratory birds increases the number of amplifying local hosts. At spatial scale, identified low - high - risk regions of public health interest, which agreed with the observed WNV cases in birds and equids. The reproductive number classified regions in the eastern parts of Germany as the main hotspot, while over time, new circulation areas were identified in the north and south. In addition, this study draws insights into the spreading pattern of WNV in Germany. Our approach, which uses the diffusion process of mosquitoes and birds, and advection for migratory birds in combination with temperature-dependent parameters, can infer the spreading speed of WNV in Germany. The simulation of the PDE model shows that the diffusion coefficients are high during the years 2023 and 2024. Due to lack of real movement data for hosts, we chose the diffusion coefficients that made the model prediction match the observed data. During the year 2023, the model prediction showed a poor fit to observed data when only resident birds were considered. However, incorporating migratory bird compartments, and capturing their directional seasonal flyways led to an improved fit for the same year. Because the pattern of observed WNV cases aligns with known migratory bird flyways, this result suggests that directed long-range bird migration could have played a significant role in transporting WNV across different locations in Germany, as also underlined by a recent study by Tóth et al. (2025). Our results imply that migratory birds may still have played an indirect role in shaping the WNV spreading pattern in Germany, even if they may not be the primary source of infection. Thus, effective surveillance and prevention efforts of WNV should also be implemented at key stopover sites, where migratory birds are likely to rest and interact with local mosquitoes. The observed data for 2023 did not validate regions in the southwest of Germany as potential hotspots, as only one single equid was confirmed in 2023 in this region. Interestingly, in 2024, the data obtained confirmed the southwestern part as a WNV circulation area, highlighting the predictive capacity of our model to accurately predict the WNV spreading pattern in Germany. The current WNV surveillance system remains patchy, compared to active surveillance (entomological, animal sentinels or blood donations) and passive surveillance systems (dead birds or symptomatic humans) that are cheaper and simpler, but are unable to detect the virus early before human cases occur. This is evidenced by that most data arises from incidental or passive detection rather than systematic monitoring (Constant et al., 2022).
Results from our spatially explicit PDE model demonstrate that temperature has a significant influence on the spreading pattern of WNV. Despite our model’s ability to simulate the spread of WNV in Germany, some additional factors can be considered to improve such PDE models in the future. For example, land use is known to influence the general distribution of mosquito populations (Watts et al., 2021), which plays a huge role in the mosquito abundance. Furthermore, we used temperature only to explain the yearly mosquito biting and mortality rates. Precipitation, wind speed, and direction also influence the general distribution and activity of mosquitoes (Ducheyne et al., 2007). The estimation of diffusion and advection coefficients by inverse modeling causes uncertainties in the results of the model, and thus, this work can be improved by considering more systematic methods for inferring diffusion and advection coefficients, such as Bayesian methods (Barajas-Solano and Tartakovsky, 2019). To further capture the random movement of hosts and mosquitoes, this PDE model can be extended to a stochastic PDE system to study more random effects that cannot be modeled by deterministic PDEs.
CRediT authorship contribution statement
Pride Duve: Designed the study, developed the model, proved the mathematical properties, computed the numerical simulations, and wrote the original draft. Renke Lühken: Supervised the study, designed the model, and reviewed the final draft. Felix Sauer: Supervised the study, designed the model, and reviewed the final draft.
Declaration of competing interests
Data availability
The cleaned WNV data for birds and equids used in this manuscript is freely available Here. Original datasets can be obtained from the World Organisation for Animal Health (2018 - 2024) (WOAH) website. The original data for human WNV cases is available from the ECDC website (European Centre for Disease Prevention and Control, 2019, 2020, 2021, 2022, 2023, 2024).
Acknowledgements
8 Appendix
@@
We write system (7) together with the corresponding initial conditions, in the Banach space of continuous functions that include the boundary, as follows (Abboubakar et al., 2023):
@@
and thus is locally Lipschitz in By [(Capasso, 1993), Theorem B.17], (Mora, 1983), and [(Smoller, 1983), Theorem 14.4], there exist a local smooth and unique solution of system (7) in We observe that system (7) can be written in the form of system (14.12) in the book (Smoller, 1983), together with initial data defined in system (14.13). By [(Smoller, 1983), Theorem (14.14)], the solutions of system (7) are always positive. ∎
@@ 8.1 Well-posedness of the model in a feasible region
We follow the approach presented in (Peng and Zhao, 2012; Wang and Dai, 2022). Using the regularity theory of parabolic PDEs (Pao, 1993), system (7) admits a unique non-negative classical solution
where represents the maximal existence time of the solution. By the strong maximum principle (Protter and Weinberger, 1984), then where are positive in
@@
we get
and are bounded for all By the positivity of and [(Peng and Zhao, 2012), Lemma (3.1)], with we conclude that there exist a positive constant that does not depend on initial data such that the solution satisfies
@@
8.2 Matrices of the PDEToolbox
while

### v2

@@ 3 Mathematical analysis of the model
We write system (6) together with the corresponding initial conditions, in the Banach space of continuous functions that include the boundary, as follows (abboubakar2023reaction):
@@
and thus is locally Lipschitz in By [(Capasso1993), Theorem B.17], (Mora1983), and [(Smoller1983), Theorem 14.4], there exist a local smooth and unique solution of system (6) in . We observe that system (6) can be written in the form of system (14.12) in the book (Smoller1983), together with initial data defined in system (14.13). By [(Smoller1983), Theorem (14.14)], the solutions of system (6) are always positive. ∎
@@ 3.1 Well-posedness of the model in a feasible region
We follow the approach presented in (Peng2012; Wang2022). Using the regularity theory of parabolic PDEs (Pao1993), system (6) admits a unique non-negative classical solution
where represents the maximal existence time of the solution. By the strong maximum principle (Protter1984), then where are positive in
@@
, we get
and are bounded for all By the positivity of and [(Peng2012), Lemma (3.1)], with we conclude that there exist a positive constant that does not depend on initial data such that the solution satisfies
@@
4 Steady states and the reproductive number
The PDE system (6) admits a non-trivial WNV-free equilibrium point such that
The uniqueness of can be verified using [(Wang2018), Lemma 2.1]. The basic reproductive number of the deterministic part of our model is computed following vandenDriessche2002, as follows:
Let be the matrix of new infections, and the matrix of transitions, then
Partial derivatives evaluated at the WNV-free equilibrium point yield:
and
The basic reproductive number is thus given by the spectral radius of which gives
The basic reproductive number can be expressed as where
with equality when which shows the contribution of migratory birds in increasing the amplifying host population. In Figure (2), the basic reproductive number is plotted for the entire country and compared to the observed data for the years 2018-2024. shows a bigger hotspot in the eastern parts of Germany, with other circulation areas observed in the Southeastern towards the Northern parts, which is in agreement with the observed data. Notably, we also included the 2025 case data to evaluate whether our risk map can predict regions where future cases may occur. Indeed, 2025 WNV cases as of November 2025 were observed in the east, north and southwest of Germany, in agreement with the high risk areas flagged by . Such a result is important as it helps identify future high-risk regions before seasonal outbreaks occur.
5 Simulation framework
Based on available WNV data in Germany, birds of the Accipitridae family were the most affected bird species in 2018 (Ziegler2022). Therefore, as also done by Mbaoma2024, we used parameters associated with northern goshawks and raptors for the resident bird, both of which fall under the Accipitridae family. Initial population densities of infectious birds and mosquitoes are estimated to be high in Halle (Saale), the main WNV hot-spot in 2018, followed by Berlin, and lastly smaller densities in Laage, Poing, and Plessa, as indicated by the 2018 WNV observed cases data (Ziegler2020). On the other hand, initial densities of susceptible hosts and mosquitoes are assumed to be spatially homogeneously distributed across the entire country. A 2-dimensional Gaussian function of the form
is applied to define the initial condition peaks. are the central coordinates defining the nodes of the infected regions, and is the standard deviation. km in Berlin and Halle Saale (main hotspot regions), and 20 km in other infected regions.
Our model is simulated from 1 March 2019 until 30 November 2024 using a daily time step. This is because no WNV activity is observed during December to February, but to account for seasonal migration patterns of migratory birds, we include the spring and autumn months. During the model simulation process, the final solution of each year is saved and imported as initial conditions for the following year. This is done to account for the inflow of new data through ongoing WNV surveillance activities. Moreover, we are able to update initial conditions for the year 2020, where a new case of WNV was confirmed in the north of Germany (Hamburg), and cases in this region may have played a role in the spreading pattern observed in the north of Germany during the subsequent years. In the process, we use inverse modeling to infer the diffusion and advection coefficients that could have contributed to the observed spreading pattern. This is done by running the model with different coefficients while visually comparing the model solution with observed data. The final coefficients are chosen based on their ability to reproduce the observed pattern, while respecting the biology of the magnitudes, i.e., The use of such an approach is mainly because available data is limited and sparse, and thus we are mainly interested in explaining the observed qualitative WNV spreading pattern.
Temperature strongly influences the activity of Cx. pipiens including their competence in transmitting WNV (Ciota2014). This motivated us to model mosquito parameters using temperature-dependent functions, namely: mosquito biting rate latency rate and mortality rate ERA5 reanalysis data on single levels from 1940 to present is downloaded from the Copernicus Climate Data Store (copcds). The data is first cleaned, and the years 2019-2024 are subsequently extracted for interpolation onto the mesh and used to compute the temperature-space dependent mosquito functions. During the simulation process, and are dynamically updated based on the temperatures at different locations in Germany. Advection representing the directed movement of migratory birds is fixed throughout the years. The spring to early summer movement is described by the vector field while the late summer to autumn is described by in km per day. To solve our PDE model, we use the Matlab PDEToolbox (MATLAB2023), which uses a finite element algorithm. The toolbox solves PDEs of the form
and in our case, Matrices and capture the diffusion coefficients and reaction terms, respectively, and they are displayed in supplementary (S4), together with the pseudo-code for the simulation procedure. The map of Germany is downloaded as a shape file from (GADM) and simplified using mapshaper (Harrower2006). The observed data indicates presence data only, i.e., no absence data is available. For this reason, we qualitatively present our results (Low - High) to interpret the model solution.
6 Model validation
To validate our model predictions, we compared simulated cases with observed data. The validation data comprises bird and equid infection events, as well as mortality, obtained from the klji. Due to the lack of a systematic WNV surveillance in Germany, many dead birds go unnoticed, while some are discovered at very advanced stages that affect PCR tests. Domesticated equids are usually easier to diagnose as owners often report their sick animals. Given that birds are the amplifying hosts, it is reasonable to make an assumption that before equid cases are reported, there is an amplifying host that was infected first and might have gone undetected. Moreover, our observed bird data is dominated by dead birds, while the equid data is dominated by infected equids. Because of this, our validation dataset compares merged WNV-related events observed in birds and equids, with the compartment of dead resident birds, given the high WNV-related deaths observed in birds of the Accipitridae family (Mbaoma2024; Ziegler2022). Firstly, we aggregate both the simulated and observed data within the GADM level-2 administrative units of Germany (N = 356) (GADM), comprising both rural districts (Landkreise) and urban districts (kreisfreie Städte). The PDE solution is extracted at the final time step, and the domain is divided into 365 sub domains corresponding to the GADM level-2 administrative units, such that For each the mean simulated number of cases is given by where is the area of the region
Observed cases are aggregated using the formula where is the indicator function equal to 1 if the observed point lies within i, and 0 otherwise, and is the number of observed points per region.
Boundaries are shrunk inwards using the polybuffer function (MATLAB2023) from the toolbox to prevent misclassifications near borders. The isinterior operator (MATLAB2023) is then applied to prevent double-counting elements located on regional borders. Thus, only points lying strictly within the interior of a polygon are counted as belonging to that region.
Lastly, the Spearman’s rank correlation coefficient (Spearman1904) is applied to quantify the degree of spatial correspondence between simulated and observed patterns in each district. Spearman’s is calculated using the formula:
where is the difference between the rank of the simulated and observed points, and is the number of administrative units. A positive indicates that locations with a high density of simulated cases correspond to locations with a high density of observed points. Conversely, a negative implies that high values in one pattern correspond to low values in the other, indicating an opposite relationship. A close to zero suggests the absence of a monotonic spatial relationship (Spearman1904). P-values are computed to assess whether the observed correlation occurred by chance, with smaller p-values (<0.05) indicating that the observed correlation is unlikely to have occurred by chance.
7 Results
The model initialized with 2018 WNV cases data shows strong visual agreement with observed patterns in Germany throughout the 2019-2024 simulation period (Figure 3). Its notable strength is the ability to identify key WNV circulation areas which are initially concentrated in the Eastern regions. The RMSD estimates indicate that WNV-mosquito infections reached an average annual dispersal range of 33.1 km in the years 2019-2022. A higher dispersal range of 46.8 km was observed in years 2023-2024 when the virus range expanded to almost the entire country. In migratory birds, an annual RMSD of 104.7 km was observed for all the years except the year 2021 where an RMSD of 93.6 km was observed. On the other hand, a constant annual dispersal range of 74.0 km was observed in resident bird infections for all the years (Table 2). The RMSD of mosquitoes remained lower than that of resident and migratory birds, while that of resident birds was lower than that of migratory birds, reflecting the greater mobility of avian hosts compared to mosquito species.
Spearman’s correlation coefficients were all statistically significant with the years 2020-2021 showing values greater than 0.5. The other years had values slightly below 0.5, a result partly attributed to the sparsity of the observed data. The corresponding p-values were all below 0.05 (Table 2, Figure 5). In 2023, the model accurately predicted an isolated case in the Northwestern region (next to the Dutch border) and another isolated case in the Southwest region between the states of Baden-Württemberg, Hessen and Rheinland-Pfalz (Figure 4(b)). This hints at the model’s ability to anticipate under-reported spread patterns mainly driven by temperature, diffusion and advection processes (Figure 3). During the same year, the model simulated without migratory birds failed to account for these isolated cases (Figure 4(a)). This shows the significance of incorporating migratory bird species and their seasonal movements in our model as this significantly improved the model fit (Figure 4(b)). In general, the spread pattern shows a progressive, wave-like expansion mainly dominated in the eastern regions initially. The wave then travels in an anti-clockwise direction to other parts of Germany. The model’s ability to explain the observed pattern suggests that incorporating temperature, diffusion and migratory processes in PDE-based WNV models significantly enhances their predictive capacity.
Year D1 D2 D3 RMSD1 RMSD2 RMSD3 P-value 2019 1.0 5.0 10.0 33.1 74.0 104.7 0.41 2020 1.0 5.0 10.0 33.1 74.0 104.7 0.53 2021 1.0 5.0 8. 0 33.1 74.0 93.6 0.51 2022 1.0 5.0 10.0 33.1 74.0 104.7 0.47 2023 2.0 5.0 10.0 46.8 74.0 104.7 0.43 2024 2.0 5.0 10.0 46.8 74.0 104.7 0.35
8 Discussions and conclusions
Understanding the WNV spreading pattern allows early detection of WNV outbreaks. The identification of circulation areas also helps early planning and resource allocation for preventive measures. At the same time, it is costly to carry out systematic surveillance programs aiming to understand the spread of WNV. Therefore, in this study, we modeled the spread of WNV in Germany using a temperature-driven PDE model with advection and diffusion terms. We proved that our model has non-negative and unique solutions that are bounded, and that the system (6) is well-posed.
In agreement with previous studies (Bhowmick2020; DBergsman2016; Mbaoma2024), at the temporal level, the basic reproductive number indicates that the presence of migratory birds increases the number of amplifying local hosts. At the spatial scale, identified low-high risk regions of public health interest, which agreed with the observed WNV cases in birds and equids, and other temperature-driven WNV models, e.g., by Mbaoma2024. The reproductive number classified regions in the eastern parts of Germany as the main hotspots, a result similar to that obtained by Bhowmick2023 and Mbaoma2024. In addition, flagged regions in the north, south and southwest of Germany as new circulation areas. Our approach explains the drivers of the WNV spreading pattern in Germany, while inferring estimates for spreading speeds. In particular, from the simulation of our PDE model, we computed the annual RMSD, which estimates the average displacement of the infection after a certain period of time. Due to a lack of real movement data for hosts, we chose diffusion coefficients that made the model prediction match the observed data.
During the year 2023, the model prediction showed a poor fit to observed data when only resident birds were considered. However, incorporating migratory bird compartments and capturing their directional seasonal flyways led to an improved fit for the same year. Observed WNV spreading pattern aligns with known migratory bird flyways, suggesting that directed long-range bird migration could play a significant role in disseminating the WNV. This observation is also supported by a recent study by Tth2025. Our results suggest that migratory birds may still have played an indirect role in shaping the WNV spreading pattern in Germany, even if they may not be the primary source of infection. Thus, effective surveillance and prevention efforts of WNV should also be implemented at key stopover sites, where migratory birds are likely to rest and interact with local mosquitoes.
The observed data for 2023 did not show any WNV cases between the northern region surrounding Hamburg and the isolated case near the Dutch border. However, our model flagged this region as a potential circulation region, a claim that was validated by the WNV data for 2024 (Figure 2 and Figure 3(f)). This highlights the predictive capacity of our model, mainly driven by the diffusion, advection and temperature. Furthermore, our model uncovers a transmission corridor for WNV that spreads from the east in an anti-clockwise direction, likely reflecting the temperature gradient that supports the long-term establishment of highly competent Cx. mosquitoes. Interestingly, the circulation areas identified in our model agree with those in the study by DiPol2022, which aimed at modeling the temperature suitability for the risk of WNV establishment in European Cx. pipiens populations. As of November 2025, WNV bird and equid cases for the year 2025 had been reported in the the Eastern, Northern and Southwestern regions, regions that were previously flagged as circulation areas in previous years (Figure 2, right).
Results from our spatially explicit PDE model demonstrate that temperature, diffusion and advection, together with resident and migratory birds have a significant influence on the spreading pattern of WNV in Germany. However, despite our model’s ability to simulate the spread of WNV in Germany, some additional factors can be considered to improve such PDE models in the future. For example, land use is known to influence the general spatio-temporal distribution of mosquito populations (Watts2021), which plays a huge role in the mosquito abundance. Furthermore, we used temperature as the only weather variable to explain the heterogeneity of our mosquito parameters. Whereas, precipitation, wind speed and direction also influence the general distribution and activity of mosquitoes (Ducheyne2007). The estimation of diffusion and advection coefficients by inverse modeling causes uncertainties in the results of the model, and thus, this work can be improved by considering more systematic methods for inferring diffusion and advection coefficients, such as Bayesian methods (BarajasSolano2019). To further capture the spatial heterogeneity in the model, another improvement would be the use of spatially varying diffusion coefficients, given that in reality, movement rates are not constant. Our PDE model can be extended to a stochastic PDE system to study more random effects that cannot be modeled by deterministic PDEs. The current WNV surveillance system remains patchy, and unable to detect the virus early before a spillover to humans occurs. This is evidenced by that most data arises from incidental or passive detection rather than systematic monitoring (Constant2022). Such a system poses a significant risk of a reporting bias in the observed data, with a potential of under-reporting of cases, prompting our results to be considered as minimum estimates, rather than the exact WNV situation.
CRediT authorship contribution statement
Pride Duve: Designed the study, developed the model, proved the mathematical properties, computed numerical simulations, and wrote the original draft. Felix Sauer: Supervised the study, designed the model, and reviewed the final draft. Renke Lühken: Supervised the study, designed the model, and reviewed the final draft.
Declaration of competing interests
Data availability
The cleaned WNV data for birds and equids used in this manuscript is freely available Here. Original datasets can be obtained from the klji website.
Acknowledgements